OpenStudy (anonymous):

How much force is needed to stop a 90-kg soccer player if he decelerates at 15 m/s²? A. 6 N B. 15 N C. 135 N D. 1350 N

OpenStudy (anonymous):

Force = mass × accelerationF = m* a m = mass = 90 kg a = acceleration = 15 m/s^2 F = force

OpenStudy (anonymous):

So what would the formula be if it were filled out?

OpenStudy (anonymous):

f=90*-15?

OpenStudy (anonymous):

\[How much force is needed \to stop a 90-kg soccer player if he decelerates at 15 m/s ^{2}\]

OpenStudy (anonymous):

Thats the question without the diamond ?s ok

OpenStudy (ipwnbunnies):

That's correct, Dragon. I guess the sign doesn't matter, since the direction is implied by the term 'decelerating'.

OpenStudy (anonymous):

@Directrix Sry to bother you again... but i need help

OpenStudy (anonymous):

@Directrix ?

Directrix (directrix):

Look at this link to the same question: http://openstudy.com/updates/5362edd4e4b0fac5fcfeb6f9

Directrix (directrix):

That is the work of @Paulodomingos

Directrix (directrix):

He wrote 1350 N. Look at his work.

Directrix (directrix):

F = m* a m = mass = 90 kg a = acceleration = 15 m/s^2 F = force @Euler271